OpenStudy (anonymous):

Rosita Can wax her car in 1 hour. when she works together with debbie, they can wax the car in 15 minutes. how long would it take debbie, working by herself, to wax the car?

OpenStudy (anonymous):

Rosita can do 1/60 job in 1 min. d=minutes it takes Debbie to do the job. Debbie can do 1/d the job in 1 min. Together they can do 1/15 the job in 1 min. 1/60 + 1/d = 1/15 d+60=4d Multiply by 60d. 3d=60 d=20 min